# DiffuBox: Refining 3D Object Detection with Point Diffusion

### v1

@@ Abstract
Ensuring robust 3D object detection and localization is crucial for many applications in robotics and autonomous driving. Recent models, however, face difficulties in maintaining high performance when applied to domains with differing sensor setups or geographic locations, often resulting in poor localization accuracy due to domain shift. To overcome this challenge, we introduce a novel diffusion-based box refinement approach. This method employs a domain-agnostic diffusion model, conditioned on the LiDAR points surrounding a coarse bounding box, to simultaneously refine the box’s location, size, and orientation. We evaluate this approach under various domain adaptation settings, and our results reveal significant improvements across different datasets, object classes and detectors.
@@ 1 Introduction
Recognizing this observation, we propose DiffuBox, a novel point diffusion model that learns the distribution of points relative to the object’s bounding box in order to refine noisy bounding box proposals from the detection models for off-the-shelf domain adaptation. Given a set of noisy bounding box proposals, DiffuBox denoises them into accurate detection boxes conditioned on the points near proposed bounding boxes. Our method naturally avoids the domain gap caused by the scale difference [47], since DiffuBox is designed to operate on object scale-invariant data, where we transform the LiDAR points around bounding box proposals into a normalized box view that is relative to the box instead of in absolute measure. This eliminates the shape priors presented in the source domain and forces the diffusion model to recover the accurate bounding box solely based on the relative position of points to the bounding box proposal, allowing for improved robustness in self-driving systems.
@@ 2 Related Work
In general, most 3D object detection methods require supervision from human-annotated data. They take 3D sensory data (e.g. LiDAR point clouds) and infer bounding boxes around 3D objects. 3D detection methods can be grouped into two categories based on the input representations: Point-based methods [32, 33, 31, 39, 40, 54, 29] that directly operate on point clouds, and grid-based methods [50, 63, 20, 38, 48, 26] that first voxelize point clouds into 3D grids and then leverage convolutional architectures. As of other supervised models, 3D detection models suffer from the decrease in performance when the distribution of data in inference is different from that in training. Our method DiffuBox is designed to reduce the domain gap for general 3D object detection, agnostic to underlining model design.
@@
Domain adaptation aims to alleviate the performance drop of 3D perception models under domain shift. [47] by Wang et al. is one of the first works studying the domain gap in 3D object detection and proposed Statistical Normalization (SN) that reduces the shape bias across domains. ST3D [52], Rote-DA [57], and ST3D++[53] propose a self-training pipeline that iteratively improve the target domain 3D detection performance with pseudo-label training and auxiliary priors. Other methods can be grouped into feature-based [49, 21, 28, 19, 37, 17] and architecture-based [12, 35, 46, 22, 36] methods. Some of them also apply data augmentation to construct and train domain-invariant representations to reduce the domain gap [19, 37, 12]. Our proposed method is orthogonal to these methods and can be applied together with these models.
@@
Recently, diffusion models [41, 9, 42, 43] have shown high-quality generative ability for image [4, 34], video [10, 8] and 3D shape [25, 59, 27] modalities. Zhou et al. [60] use diffusion models with a point-voxel representation for shape generation and point-cloud completion. LION [59] uses a hierarchical VAE mapped to a latent space and trains diffusion models on latent encodings to generate point clouds. In perception tasks, Chen et al. [3] and Zhou et al. [62] propose diffusion-based object detection frameworks. Kim et al. [18] proposes a diffusion-based module to enhance the proposal refinement stage of two-stage object detectors. Differently, our approach focuses on leveraging diffusion for post-processing in a detector-agnostic manner that shows superior performance over previous approaches.
@@ 3.3 Bounding Box Refinement via Diffusion
Figure 2 illustrates our hypothesis that the good localization of a bounding box $\boldsymbol{b}$ is closely correlated to its corresponding $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}$ forming a “standard” point distribution, a.k.a. shape. In return, improving the shape of $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}$ will also lead to better localization of $\boldsymbol{b}$. Since our ultimate goal is to optimize the bounding box $\boldsymbol{b}\rightarrow\hat{\boldsymbol{b}}$, we propose to use a point diffusion model to learn to “denoise” $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}\rightarrow\boldsymbol{P}_{\hat{\boldsymbol{b}}}^{\text{NBV}}$.
Specifically, we begin by discussing the training of the diffusion model to learn the probabilistic flow for each point to a good box in Section 3.3.1. Then, we go into how our method refines the bounding box by computing the improvement step relative to the learned probabilistic flow in Section 3.3.2. Finally, we go into how we leverage the shape guidance to embed heuristics into our training procedure in Section 3.3.3.
@@ 4.1 Experimental Setup
We consider three datasets: The KITTI dataset [6], the Lyft Level 5 Perception dataset [16], and the Ithaca365 dataset [5]. For KITTI, we follow the official splits. For Lyft, we follow various existing work [56, 57, 24] and use the splits separated by geographical locations, consisting of 11,873 point clouds for training and 4,901 for testing. For Ithaca365, we utilize the annotated point clouds with 4,445 for training and 1,644 for testing.
@@
We use the implementation and configurations from OpenPCDet [44] for detectors, and [14]’s implementation for diffusion models. We use context limit of 4x the bounding box size, and shape weight 0.1 for cars and pedestrians and 0.01 for cyclists as cyclists have more shape variation. More details can be found in the supplementary.
@@ 4.2 Experimental Results
We present the results for KITTI $\rightarrow$ Lyft cars in Table 1 (mAP@IoU 0.7) and Table S8 (nuScenes TP metrics), and the results for KITTI $\rightarrow$ Ithaca365 cars in Table 2 (mAP@IoU 0.7) and Table S9 (nuScenes TP metrics). We use PointRCNN [39] detectors; evaluations with other detectors can be found in the section below. DiffuBox consistently attains significant performance gain across different domain adaptation methods and datasets. The improvement of DiffuBox is especially significant for near-range and middle-range detections. Notably, for KITTI $\rightarrow$ Lyft cars, DiffuBox applied upon the Direct outputs is able to attain comparable performance with domain adaptation methods that require training such as ST3D. We hypothesize that this is because there are more LiDAR points for near-range and middle-range objects, which allows DiffuBox to better correct the detections.
Table 1: mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft cars. Higher is Better. DiffuBox leads to improvement in almost all cases, with especially significant gain for the Direct and OT detections.
@@
Table 2: mAP@IoU 0.7 for KITTI $\rightarrow$ Ithaca365 cars. Higher is Better. DiffuBox leads to significant improvement upon different adaptation methods.
@@
To further demonstrate the robustness and versatility of DiffuBox, we present DiffuBox’s performance for refining predictions from other detectors. We consider PointPillar [20], SECOND [50] and PV-RCNN [38] trained on KITTI. For PV-RCNN, we further consider its combination with OT and SN. We report the mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft cars in Table 3. Results show that DiffuBox consistently improves the predictions from different detectors.
Table 3: mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft cars with other detectors. Higher is better. DiffuBox consistently improves the detections from different detectors.
@@
Table 4: mAP@IoU 0.5 for KITTI $\rightarrow$ Lyft pedestrians and cyclists. Higher is better. DiffuBox attains consistent improvement across object classes and is able to improve upon other domain adaptation methods.
@@ 4.3 Qualitative Results
Figure 3 visualizes four scenes from Lyft. We compare the ground truth bounding boxes (green), the detections directly obtained from a PointRCNN trained on KITTI (yellow), and the refined detections using DiffuBox (blue). The out-of-domain PointRCNN produces reasonable results, but occasionally it produces false positives or boxes with incorrect shapes or alignment. DiffuBox effectively moves the incorrect boxes towards having better location, shape and alignment. Also observe that for the already accurate boxes, DiffuBox makes little change to them.
Figure 3: Illustration of 3D object detection on the Lyft/Ithaca365 dataset before and after DiffuBox’s refinement. We visualize detections from an out-of-domain PointRCNN on four scenes from each dataset. We color the ground truth boxes in green, the detector outputs in yellow, and DiffuBox’s refinements in blue. The out-of-domain detector sometimes produces false positives or boxes with incorrect shape or alignment. DiffuBox effectively improves the wrong or inaccurate boxes, while making little change to the accurate boxes.
@@ 4.4 DiffuBox Extension: Detector Retraining
We provide the KITTI $\rightarrow$ Lyft results in Table 5. We compare the performance of (1) directly applying the KITTI detector, (2) retraining for one round with the KITTI detector’s predictions as labels, and (3) retraining for one round with the KITTI detector’s predictions after DiffuBox’s refinement as labels. Retraining is performed based on the KITTI detector’s predictions on the Lyft training split, and the evaluation is conducted over Lyft’s testing split. Retraining using the KITTI detector’s predictions directly only provides limited improvement over directly applying the KITTI detector, while retraining with DiffuBox’s refinement leads to significant improvement in performance.
Table 5: Retraining Performance (mAP@IoU 0.7) for KITTI $\rightarrow$ Lyft. Higher is better. Retraining with DiffuBox’s refined detections attains significant performance improvement.
@@ 4.5 Ablation Studies and Analysis
We conduct ablation study on the context limit, and consider ranges of 2x, 4x and 6x the box size. We evaluate under the setting of no adaptation (Direct) for KITTI $\rightarrow$ Lyft cars. We compare the mAP@IoU 0.7 in Table 6. All three context limits lead to significant performance improvement. Larger limit attains more gain, and the gain saturates at around the 4x limit.
@@
We visualize the comparison of IoU of the bounding boxes with the ground truth bounding boxes both before and after using DiffuBox in Figure 4. The IoU of predictions after using our method (in blue) improves significantly over before refinement (in yellow). This suggests that a majority of our refinement is in improving the bounding boxes’ shape and alignment to fit into the new domain, thus resulting in higher IoU values.
@@ 5 Discussion and Future Work
In this work, we propose DiffuBox, a diffusion-based approach that refines bounding boxes for better domain adaptation. While DiffuBox effectively improves the existing bounding boxes, one limitation is that DiffuBox currently does not consider false negatives. This could potentially be addressed through further distilling the detectors by DiffuBox’s refined boxes, or by incorporating exploration strategies to capture potential false negatives. Alternatively, a view that we did not discuss in this work, but is of potential interest to the field, is the use of DiffuBox for automatic label refinement. This can be useful for correcting slightly mis-aligned boxes, or labels across sensors that may have slightly unsynchronized sensors. We leave this discussion for further work, and will provide code and model checkpoints for this use case. With our work, we do not foresee any negative societal impacts and hope the field continues to develop such label refinement methods for 3D object detection.
@@
We use the implementation and configuration from OpenPCDet [44]. We use [14]’s implementation for the diffusion model and follow their noise schedule $\sigma_{max}=80$. During inference, we denoise each box for 16 steps, using a noise level between $[10,80]$ linear in the detector confidence; higher confidence boxes begin from smaller noise. For each bounding box, we consider points falling within 4x the bounding box size. Additionally, inspired by OT and SN, we slightly regularize the bounding box shape with the average size of the target domain, with shape weight 0.1. We use NVIDIA A6000 for all of our experiments.
@@
In the diffusion process, we follow [13] and use noise level distribution $\ln\sigma\sim\mathcal{N}\left(-1.2,{1.2}^{2}\right)$, ODE schedule $\sigma(t)=t$, and $2^{\text{nd}}$ order Heun solver. The denoiser transformer model contains 12 self-attention layers with hidden size 1024, each layer has 2048 intermediate dimensions and 8 heads. The diffusion model is trained with batch size 128, learning rate 0.0001, for 100k steps. A visualization of our Diffusion model architecture can be seen in Figure S5.
@@
S2.1 nuScenes True Positive Metrics
We report the nuScenes TP metrics for KITTI $\rightarrow$ Lyft cars in Table S8, and KITTI $\rightarrow$ Ithaca365 in Table S9. All results are based on PointRCNN detectors. DiffuBox effectively reduces all three types of errors, with especially significant improvement for translation and scale errors.
Table S8: nuScenes TP metrics for KITTI $\rightarrow$ Lyft cars. Lower is better. DiffuBox reduces all three types of errors, with especially significant improvement for translation and scale errors.
@@
Table S9: nuScenes TP metrics for KITTI $\rightarrow$ Ithaca365 cars. Lower is better.
@@
S2.2 Performance on Ithaca365 Pedestrians
We report the performance on Ithaca365 pedestrians using PointRCNN detectors in Table S10. DiffuBox attains significant improvement upon the detections.
Table S10: mAP@IoU 0.5 for KITTI $\rightarrow$ Ithaca365 pedestrians. Higher is better. DiffuBox consistently attains significant improvement.

### v2

@@ Abstract
Ensuring robust 3D object detection and localization is crucial for many applications in robotics and autonomous driving. Recent models, however, face difficulties in maintaining high performance when applied to domains with differing sensor setups or geographic locations, often resulting in poor localization accuracy due to domain shift. To overcome this challenge, we introduce a novel diffusion-based box refinement approach. This method employs a domain-agnostic diffusion model, conditioned on the LiDAR points surrounding a coarse bounding box, to simultaneously refine the box’s location, size, and orientation. We evaluate this approach under various domain adaptation settings, and our results reveal significant improvements across different datasets, object classes and detectors. Our PyTorch implementation is available at https://github.com/cxy1997/DiffuBox.
@@ 1 Introduction
Recognizing this observation, we propose DiffuBox, a novel point diffusion model that learns the distribution of points relative to the object’s bounding box in order to refine noisy bounding box proposals from the detection models for off-the-shelf domain adaptation. Given a set of noisy bounding box proposals, DiffuBox denoises them into accurate detection boxes conditioned on the points near proposed bounding boxes. Our method naturally avoids the domain gap caused by the scale difference [48], since DiffuBox is designed to operate on object scale-invariant data, where we transform the LiDAR points around bounding box proposals into a normalized box view that is relative to the box instead of in absolute measure. This eliminates the size priors presented in the source domain and forces the diffusion model to recover the accurate bounding box solely based on the relative position of points to the bounding box proposal, allowing for improved robustness in self-driving systems.
@@ 2 Related Work
In general, most 3D object detection methods require supervision from human-annotated data. They take 3D sensory data (e.g. LiDAR point clouds) and infer bounding boxes around 3D objects. 3D detection methods can be grouped into two categories based on the input representations: Point-based methods [32, 33, 31, 39, 40, 55, 29] that directly operate on point clouds, and grid-based methods [51, 65, 20, 38, 49, 26] that first voxelize point clouds into 3D grids and then leverage convolutional architectures. Like other supervised models, 3D detection models suffer from decreased performance when the data distribution during inference differs from that during training. Our method DiffuBox is designed to reduce the domain gap for general 3D object detection, agnostic to underlining model design.
@@
Domain adaptation aims to alleviate the performance drop of 3D perception models under domain shift. [48] by Wang et al. is one of the first works studying the domain gap in 3D object detection and proposes Statistical Normalization (SN) that reduces the shape bias across domains. ST3D [53], Rote-DA [59], and ST3D++ [54] propose a self-training pipeline that iteratively improve the target domain 3D detection performance with pseudo-label training and auxiliary priors. Other methods can be grouped into feature-based [50, 21, 28, 19, 37, 17] and architecture-based [12, 35, 46, 22, 36] methods. Some of them also apply data augmentation to construct and train domain-invariant representations to reduce the domain gap [19, 37, 12]. Our proposed method is orthogonal to these methods and can be applied together with these models.
@@
Recently, diffusion models [41, 9, 42, 43] have shown high-quality generative ability for image [4, 34], video [10, 8] and 3D shape [25, 61, 27] modalities. Zhou et al. [62] uses diffusion models with a point-voxel representation for shape generation and point-cloud completion. LION [61] uses a hierarchical VAE mapped to a latent space and trains diffusion models on latent encodings to generate point clouds. In perception tasks, Chen et al. [3] and Zhou et al. [64] propose diffusion-based object detection frameworks. Kim et al. [18] proposes a diffusion-based module to enhance the proposal refinement stage of two-stage object detectors. Unlike these approaches, our work focuses on leveraging diffusion for post-processing in a detector-agnostic manner that shows superior performance over previous methods.
@@ 3.3 Bounding Box Refinement via Diffusion
Figure 2 illustrates our hypothesis that the good localization of a bounding box $\boldsymbol{b}$ is closely correlated to its corresponding $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}$ forming a “standard” point distribution, a.k.a. shape. Thus, improving the shape of $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}$ will also lead to better localization of $\boldsymbol{b}$. Since our ultimate goal is to optimize the bounding box $\boldsymbol{b}\rightarrow\hat{\boldsymbol{b}}$, we propose to use a point diffusion model to learn to “denoise” $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}\rightarrow\boldsymbol{P}_{\hat{\boldsymbol{b}}}^{\text{NBV}}$.
Specifically, we begin by discussing the training of the diffusion model to learn the probabilistic flow for each point to a good box in Section 3.3.1. Then, we go into how our method refines the bounding box by computing the improvement step relative to the learned probabilistic flow in Section 3.3.2. Finally, we go into how we leverage the shape guidance to embed heuristics into our training procedure in Section 3.3.3. To enhance clarity, we additionally provide an algorithmic description of DiffuBox’s training and inference workflow in Appendix S1.
@@ 4.1 Experimental Setup
We primarily consider three datasets: The KITTI dataset [6], the Lyft Level 5 Perception dataset [16], and the Ithaca365 dataset [5]. For KITTI, we follow the official splits. For Lyft, we follow various existing works [58, 59, 24] and use the splits separated by geographical locations, consisting of 11,873 point clouds for training and 4,901 for testing. For Ithaca365, we utilize the annotated point clouds with 4,445 for training and 1,644 for testing. Additionally, we include experiments with the nuScenes dataset in the supplementary to evaluate DiffuBox’s performance on larger-scale and more diverse data.
@@
We use the implementation and configurations from OpenPCDet [44] for detectors, and [14]’s implementation for diffusion models. We set the context limit to 4x the bounding box size. We use shape weight 0.1 for cars and pedestrians, and 0.01 for cyclists as cyclists have more shape variation. More details can be found in the supplementary.
@@ 4.2 Experimental Results
We present the results for KITTI $\rightarrow$ Lyft cars in Table 1 (mAP@IoU 0.7) and Table S8 (nuScenes TP metrics), and the results for KITTI $\rightarrow$ Ithaca365 cars in Table 2 (mAP@IoU 0.7) and Table S9 (nuScenes TP metrics). We additionally include KITTI $\rightarrow$ nuScenes results in Table S10 in the supplementary. We use PointRCNN [39] detectors; evaluations with other detectors can be found in the section below. DiffuBox consistently attains significant performance gain across different domain adaptation methods and datasets. The improvement of DiffuBox is especially significant for near-range and middle-range detections. Notably, for KITTI $\rightarrow$ Lyft cars, DiffuBox applied upon the Direct outputs is able to attain comparable performance with domain adaptation methods that require training such as ST3D. We hypothesize that this is because there are more LiDAR points for near-range and middle-range objects, which allows DiffuBox to better correct the detections.
Table 1: mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft (cars). Higher is Better. DiffuBox leads to improvement in almost all cases, with especially significant gain for the Direct and OT detections.
@@
Table 2: mAP@IoU 0.7 for KITTI $\rightarrow$ Ithaca365 (cars). DiffuBox leads to significant improvement upon different adaptation methods.
@@
To further demonstrate the robustness and versatility of DiffuBox, we present DiffuBox’s performance for refining predictions from other detectors. We consider PointPillar [20], SECOND [51], PV-RCNN [38], CenterPoint [56] and DSVT [47] trained on KITTI. We report the mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft cars in Table 3. Results show that DiffuBox consistently improves the predictions from different detectors.
Table 3: mAP@IoU 0.7 for KITTI $\rightarrow$ Lyft (cars) with other detectors. DiffuBox consistently improves the detections from different detectors.
@@
CenterPoint (Direct) 74.91 36.64 2.47 36.23 28.63 4.05 0.15 10.29
CenterPoint (Direct)+DiffuBox 90.25 58.65 6.95 51.38 71.02 34.91 1.48 34.40
CenterPoint (OT) 82.81 51.02 6.26 46.10 25.34 12.35 0.52 13.34
CenterPoint (OT)+DiffuBox 91.79 59.24 7.45 51.88 63.01 32.53 1.24 30.95
DSVT (Direct) 68.93 47.49 11.32 41.77 33.72 11.74 1.42 15.67
DSVT (Direct)+DiffuBox 89.01 63.41 17.50 55.27 65.22 36.31 5.02 35.61
DSVT (OT) 71.85 42.93 13.18 43.05 15.66 4.47 0.31 8.06
DSVT (OT)+DiffuBox 90.21 63.50 17.91 55.61 56.84 31.19 4.63 31.12
@@
Table 4: mAP@IoU 0.5 for KITTI $\rightarrow$ Lyft (pedestrians/cyclists). DiffuBox attains consistent improvement across object classes and is able to improve upon other domain adaptation methods.
@@ 4.3 Qualitative Results
Figure 3 visualizes four scenes from the Lyft and Ithaca365 datasets. We compare the ground truth bounding boxes (green), the detections directly obtained from a PointRCNN trained on KITTI (yellow), and the refined detections using DiffuBox (blue). The out-of-domain PointRCNN produces reasonable results, but occasionally it produces false positives or boxes with incorrect shapes or alignment. DiffuBox effectively moves the incorrect boxes towards having better location, shape and alignment. Also observe that for the already accurate boxes, DiffuBox makes little change to them.
Figure 3: Illustration of 3D object detection on Lyft/Ithaca365 before and after DiffuBox’s refinement. We visualize detections from an out-of-domain PointRCNN on four scenes from each dataset. We color the ground truth boxes in green, the detector outputs in yellow, and DiffuBox’s refinements in blue. The out-of-domain detector sometimes produces false positives or boxes with incorrect shape or alignment. DiffuBox effectively improves the wrong or inaccurate boxes, while making little change to the accurate boxes.
@@ 4.4 DiffuBox Extension: Detector Retraining
We provide the KITTI $\rightarrow$ Lyft results in Table 5. We compare the performance of (1) directly applying the KITTI detector, (2) retraining for one round with the KITTI detector’s predictions as labels, and (3) retraining for one round with the KITTI detector’s predictions after DiffuBox’s refinement as labels. Retraining is performed based on the KITTI detector’s predictions on the Lyft training split, and the evaluation is conducted over Lyft’s testing split. Retraining using the KITTI detector’s predictions directly only provides limited improvement over directly applying the KITTI detector, while retraining with DiffuBox’s refinement leads to significant improvement.
Table 5: Retraining Performance (mAP@IoU 0.7) for KITTI $\rightarrow$ Lyft. Retraining with DiffuBox’s refined detections attains significant performance improvement.
@@ 4.5 Ablation Studies and Analysis
We conduct ablation study on the context limit, and consider ranges of 2x, 4x and 6x the box size. We evaluate under the setting of no adaptation (Direct) for KITTI $\rightarrow$ Lyft cars, and compare the mAP@IoU 0.7 in Table 6. All three context limits lead to significant performance improvement. Larger limit attains more gain, and the gain saturates at around the 4x limit.
@@
Denoising Steps.
We perform ablation on the number of denoising steps used, and present the results in Figure 5. In general, a majority of the performance is already reached using 8 diffusion steps, and it saturates around using 14 steps.
Figure 5: mAP vs. Number of Diffusion Steps. We report the BEV (left) and 3D (right) mAP @ IoU 0.7 for the setting of KITTI $\rightarrow$ Lyft Cars and PointRCNN detector.
@@
We visualize the comparison of IoU of the bounding boxes with the ground truth bounding boxes both before and after using DiffuBox in Figure 4. The IoU of predictions after DiffuBox refinement (in blue) improves significantly over those before refinement (in yellow). This suggests that a majority of our refinement is in improving the bounding boxes’ shape and alignment to fit into the new domain, thus resulting in higher IoU values.
Recall Analysis.
We perform analysis on DiffuBox’s effect on detection recall in Figure 6. As DiffuBox improves IoU for mislocalized detections, it reduces false negatives that arise from match IoU being lower than the threshold. The improvement is observed for objects across different sizes.
(a) Recall vs. IoU
(b) Recall vs. Object volume
Figure 6: Recall improvement with DiffuBox. We report recall on the car class (KITTI $\rightarrow$ Lyft, PointRCNN) before and after refinement with DiffuBox.
@@ 5 Discussion and Future Work
In this work, we propose DiffuBox, a diffusion-based approach that refines bounding boxes for better domain adaptation. While DiffuBox effectively improves the existing bounding boxes, one limitation is that DiffuBox currently does not consider false negatives that are due to completely missed detections. This could potentially be addressed through further distilling the detectors by DiffuBox’s refined boxes, or by incorporating exploration strategies to capture possibly missed objects. Alternatively, a view that we did not discuss in this work, but is of potential interest to the field, is the use of DiffuBox for automatic label refinement. This can be useful for correcting slightly mis-aligned boxes, or labels across sensors that may have slightly unsynchronized sensors. We leave this discussion for further work, and will provide code and model checkpoints for this use case. With our work, we do not foresee any negative societal impacts and hope the field continues to develop such label refinement methods for 3D object detection.
Acknowledgments and Disclosure of Funding
This research is supported by grants from the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310). The Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, Nvidia, and the New York Presbyterian Hospital.
@@
Appendix S1 Algorithmic Description of DiffuBox
Below we provide an algorithmic description of the training and inference workflow of DiffuBox.
Algorithm 1 DiffuBox Training
1:Train set $\mathcal{D}_{\text{train}}$, noise distribution $p_{\text{train}}$
2:Diffusion model $F_{\theta}$ to approximate $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},\sigma)$ of bounding box over points $\boldsymbol{x}$, noise $\sigma$
3:repeat
4:     Sample point cloud $\boldsymbol{P}$, true bounding box $\boldsymbol{b^{*}}$ where $\boldsymbol{P},\boldsymbol{b^{*}}\sim\mathcal{D}_{\text{train}}$
5:     Sample noise level $\sigma\sim p_{\text{train}}$
6:     Compute $\boldsymbol{P}_{\boldsymbol{b^{*}}}^{\text{NBV}}$ from Eq. 2
7:     Update $F_{\theta}$ with diffusion objective Eq. 4
8:until convergence
Algorithm 2 DiffuBox Inference
1:Set of initial bounding boxes $\{\boldsymbol{b}\}_{\text{scene}}$, point cloud $\boldsymbol{P}$, $F_{\theta}$ approximating $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},\sigma)$
2:Set of refined bounding boxes $\{\hat{\boldsymbol{b}}\}_{\text{scene}}$
3:for each bounding box $\boldsymbol{b}$ in $\{\boldsymbol{b}\}_{\text{scene}}$ do
4:     for denoising steps $t=1,\cdots,T$ do
5:         Obtain $\sigma(t)$ from denoising schedule
6:         Compute $\boldsymbol{P}_{\boldsymbol{b}}^{\text{NBV}}$ from Eq. 2
7:         Compute $\text{d}\boldsymbol{b}$ with $\sigma(t)$ and $F_{\theta}$ from Eq. 8
8:         $\boldsymbol{b}\leftarrow\boldsymbol{b}+\text{d}\boldsymbol{b}$
9:     end for
10:     $\hat{\boldsymbol{b}}\leftarrow\boldsymbol{b}$
11:end for
12:$\{\hat{\boldsymbol{b}}\}_{\text{scene}}\leftarrow\text{NMS}(\{\hat{\boldsymbol{b}}\}_{\text{scene}})$
@@
For detectors, we use the implementation and configurations from OpenPCDet [44]. For diffusion models, we use [14]’s implementation and follow their noise schedule $\sigma_{max}=80$. During inference, we denoise each box for 14 steps, starting from a noise level between $[10,80]$ linear in the detector confidence; higher confidence boxes begin from smaller noise. For each bounding box, we consider the points that fall within the range of four times its size. Additionally, inspired by OT and SN, we slightly regularize the bounding box shape with the average size of the target domain, with shape weight 0.1 for cars and pedestrians, and 0.01 for cyclists. We use NVIDIA A6000 for all of our experiments.
@@
In the diffusion process, we follow [13] and use noise level distribution $\ln\sigma\sim\mathcal{N}\left(-1.2,{1.2}^{2}\right)$, ODE schedule $\sigma(t)=t$, and $2^{\text{nd}}$ order Heun solver. The denoiser transformer model contains 12 self-attention layers with hidden size 1024; each layer has 2048 intermediate dimensions and 8 heads. The diffusion model is trained with batch size 128 and learning rate 0.0001 for 100k steps. A visualization of our Diffusion model architecture can be found in Figure S7.
@@
S3.1 nuScenes True Positive Metrics
We report the nuScenes TP metrics for KITTI $\rightarrow$ Lyft cars in Table S8, and those for KITTI $\rightarrow$ Ithaca365 in Table S9. All results are based on PointRCNN detectors. DiffuBox effectively reduces all three types of errors, with especially significant improvement for translation and scale errors.
Table S8: nuScenes TP metrics for KITTI $\rightarrow$ Lyft (cars). Lower is better. DiffuBox reduces all three types of errors, with especially significant improvement for translation and scale errors.
@@
Table S9: nuScenes TP metrics for KITTI $\rightarrow$ Ithaca365 (cars). Lower is better.
@@
S3.2 Additional Experimental Results
We report the performance on KITTI $\rightarrow$ nuScenes cars in Table S10, and KITTI $\rightarrow$ Ithaca365 pedestrians in Table S11. Both sets of experiments are based on PointRCNN detectors. For nuScenes, we evaluate for up to 50m based on the point cloud range. DiffuBox attains significant improvement upon the detections in both settings.
Table S10: mAP@IoU 0.7 for KITTI $\rightarrow$ nuScenes (cars). DiffuBox consistently attains significant improvement.
Method BEV $\uparrow$ 3D $\uparrow$
 0-30m  30-50m  0-50m  0-30m  30-50m  0-50m
Direct 44.78 0.70 15.86 14.82 0.00 4.66
Direct+DiffuBox 58.07 1.06 20.70 22.77 0.00 7.40
Table S11: mAP@IoU 0.5 for KITTI $\rightarrow$ Ithaca365 (pedestrians). DiffuBox consistently attains significant improvement.